OpenStudy (anonymous):

A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = distance over time; 1 mile = 1609 meters.

OpenStudy (muzzack):

so we can do 1609 • 70 --------- 3 h

OpenStudy (mathstudent55):

speed = distance over time speed = 70 miles / 3 h \(speed= \dfrac{70~miles}{3~hours} = \dfrac{70}{3} mph = 23 \dfrac{1}{3} mph\approx 23.3 mph\)